OpenStudy (anonymous):

What is f(-5) if f(x)= |2x-1|+10

OpenStudy (anonymous):

21 19 -1 -21

OpenStudy (nubeer):

replace x with -5 and solve the equation

OpenStudy (anonymous):

So it would be -1?

OpenStudy (anonymous):

[f(5) = |2(5) + 1| = 11 f(21) = |2(21) +1| = 23 f(19)= |2(19) +1| = 39 f(-1) = |2(-1) + 1| = -1 f(-21) = |2(-21) + 1| = -21

OpenStudy (anonymous):

21 would be answer

OpenStudy (anonymous):

I meant -5 f(x) is 10

OpenStudy (anonymous):

Thank you all

OpenStudy (anonymous):

Welcome